Find the intervals in which f (x) =
|x-2|
x2
is strictly increasing and strictly decreasin

Grade:6

2 Answers

ruchi yadav
10 years ago
F(x) = (x-2)x2 ; x>=2; Strictly increasing

(2-x)x2 ; x<2; Strictly increasing



Thank You
Ruchi
IIT Delhi
Askiitians Faculty
ruchi yadav
10 years ago
F(x) = (x-2)x2 ; x>=2; Strictly increasing

(2-x)x2 ; x<2; Strictly increasing



Thank You
Ruchi
IIT Delhi
Askiitians Faculty